A meeting of the elders was held in a village in the early 1400's. In the meeting, the date of a future ceremony was first set, and the attendees were instructed to spread the date to their fellow villagers. In the table below, t represents the number of days that had passed since the meeting, and V represents the number of villagers that had acquired the date of the ceremony. 28 1225 24 1195 21 1150 16 945 460 190 125 Use the data to construct a scatter plot, then complete the following. 1) Using the scatter plot and the context of the problem, which of the following best describes the patterm? OA. Quadratic (y-ax2+bx+ OB. Exponential (y=a-bx) Logistic yhy - bx 1 +a e 2) Using your calculator and the best of the four methods above, find the model that approximates the number of villagers that had acquired the date of the ceremony t days after the meeting. Use that result to answer the following questions. ROUND TO THE NEAREST PERSON 0 D. Linear (y = mx + b) Approximately how many villagers were at the meeting of the elders? About 125 villagers were at the meeting
Approximately how many villagers were at the meeting of the elders? About 125 villagers were at the meeting. According to the model, about villagers will eventually know the date of the ceremony. How fast was the date of the ceremony spreading exactly 15 days after the meeting? It was spreading at a rate of aboutvillagers per day.
Answer #1

we have plotted the graph by given data and had find out the slope at x=15 because it will only tell the rate of spreading per village.

slope =Y/X=881/15=59

therefore it was spreading at a rate of 59 villager per day1105 965 825 X 15.03 Y-881 685 545 405 265 17. 20.222.0 9.8 25.4 4.6 7.2 15 12.4